OpenStudy (courtstacks98):

Prove the identity: (cos x + cos y)^2 + (sin x + sin y)^2=2+2cos(x+y)

Nnesha (nnesha):

expand both parentheses (cos x +cos y)^2 = ??? )Foil) (sin x +siny)^2 = ?

Nnesha (nnesha):

and after that you have to use the identity \[\rm \cos^2 \theta + \sin^2 \theta = 1\]

OpenStudy (universetoast):

Here's how you can start it: Break it up into familiar functions. (Cosx+Cosy)to the 2 is equavalent to (Cosx+Cosy)(Cosx+Cosy) You foil this and end up with: Cos^2x+Cos^y+2cosxcosy You do the same thing to the other foilable function: (Sinx+Siny) And then solve

OpenStudy (courtstacks98):

So, when I foil I do, (cos x + cos y) (cos x+ cos y) and get 4 cos x + cos(x+y) +2 cos y?

OpenStudy (courtstacks98):

oh okay nevermind to what I just said lol

Nnesha (nnesha):

 it would be easy if you draw a box like this